We’ve been busy processing up the RIVET I data from the New River Inlet, which we will be posting to our ftp site soon. We’ve also been working up the airborne observations (in collaboration with Luc Lenain & the Melville lab). The video below shows a visualization of the May 7th dye release.
Animation of repeated airborne passes of the dye plume imaged with the hyperspectral imager with New River Inlet bathymetry contours. Data is from May 7th 2012. The dye signal is constructued by differencing light in two different wavenumber bands.
The full movie file which has much more detail can be downloaded here
Here are the movies from the sonars on our quadpods. I added two new ones with 2-axis pencil beam rotary sonar imagery overlaid on top of the sidescan data to show some actual elevations of the bedforms in a limited region under the frame.
The waves at New River Inlet are supposed to be below 50 cm later this week. I looked at the waves at the buoy every morning since February. They were bigger the 4 weeks the RIVET sensors were in the water than at any time since February! There were two pre-June tropical storms. When a few of us came to New River to poke around in October 2010, we drove thru Tropical Storm Nicole (21″ of rain in 24 hours). What are the odds…
Big waves made in-water work difficult, but also provide an opportunity to see the effect of waves on inlet processes, increasing the coolness of everyone’s data sets.
No one is left besides Britt and me (waiting for trucks to take our gear to Duck) (and of course, Greg Schmidt). Quiet. I think the Tiki Bar is considering closing until late June owing to lack-of-business. There are lots of empty tables at Rick’s, and the cul-de-sac at the north end of New River Inlet road is deserted. Not even a LARC in sight.
REMUS operations starting ~1.5 hours before max ebb were conducted during the 5/19 and 5/20 dye releases. Conditions were rough during the 5/19 mission outside of the inlet and no dye was visible during REMUS ops. However comparison plots of river Colored Dissolved Organic Matter (CDOM) versus Rhodamine concentrations show that the dye was present, just at low concentrations (~0.5 ppb). The mission was designed to track the advecting river plume from the inlet mouth to the river plume extent observed from earlier missions (~ 1.2 km offshore). Mission time was ~3.5 hrs.
The same mission was performed during the dye release on 5/20/2012. Significantly higher concentrations of dye (6 – 10 ppb) were advected outside of the inlet and were observed in the evolving river plume.
Sorry for the late post, but… Thursday May 17 was the last day of SIO drifter releases. Again weather (rain and potential T-storms) cut the releases short and we only did two releases. Both releases were during ebb tide and were generally similar (thus only the first is examined in detail here) but novel compared to previous releases during ebb tide. We’ll get to the novelty shortly… For both releases all 35 drifters were released upstream of the inlets “S” turn, however, downstream of previous ebb releases which were released near the ICW. We hoped that by spreading them out at this location (where the inlet is wider) we might avoid the mass convergence that we had seen in the past. The mass convergence still occurred!
Fig.1 Tracks of drifters during ebb tide. Notice the convergence through the “S”. During this release, the exit of the drifters was different than previous ebb releases in which drifters were released upstream of the “S”. Notice that some drifters (blue) are exiting the old channel while others (red) are exiting the new channel. Previously, converged drifters exited either the “old” or “new” channel.
Fig 2. Blue drifters quickly exiting the “old” channel and red drifters sampling the inlet width: the last two (21,26) are completely in the “new” channel and the others are spread over the shoal between new and old channels.
Fig. 3. The exit location of the drifter is spread out entirely over the inlet mouth. Many drifters exit through the “old” channel heading north, many over the shoal between “new” and “old”, and 2 exit the “new” channel and head south. This spreading of the exit had not been seen in previous releases. What dictates (tide phase, tide magnitude, wind, waves?) the final drifter location is not known.
Yesterday we completed the final set of SWIFT runs, measuring surface turbulence, waves, winds, and currents through the inlet. Below is the composite of all tracks from 26 Apr to 21 May, color scaled by turbulent dissipation rate. Across all conditions, there is consistent enhanced dissipation at the channel outlets (old and new), as well as over the shoals. Other, more localized features, such as enhanced turbulence crossing fronts (and dye plumes!) are evident in the daily runs. Preliminary data products and summaries are at http://herschel.apl.washington.edu/darla/SWIFT/
The dye release on Sat 19 May was interesting. 30 gallons of dye were released for 5 hours starting at 0430 (AM!) in approximately 10 m water depth a few meters (3-4 m) above the bed between buoy #6 and #8 in the channel Dye went up the inlet towards the ICW on the flood. Dye that had reached the surface could also be observed to enter the marsh at the start of the S-turn. By the start of the ebb, the marsh was really pink with Rhodamine concentrations of approximately 10 ppb. As sampling occured on the ebb tide, concentrations of the order of 1-2 ppb were observed coming out of the inlet. This had lead some specularion that the marsh stored a signficant fraction of the dye. But that is not the case.
First the dye was released in the lower water column in 10 m depth. Only the top 0.5 m at most fo the water column could make it into the marsh. Most of the dye had to make it’s way up the inlet. Second, dye was input at 5 mL/sec at a concentration of 2e+8 ppb. 1 mL = 1 cm^3 = 1e-6 m^3. So the pumped flux of dye into the inlet is about Q=1000 ppb m^3/s.
To balance this coming out of the inlet, one needs a water velocity V, a depth h, and channel width L, and a concentration D – Assuming a rectangular channel and a uniform velocity and dye concentration. Then the flux is Q = L*h*V*D.
Taking representative values for V=1 m/s, h= 5 m, L=200 m this implies that to balance the input flux one needs….. drum roll…… D = 1 ppb. Which is the order of magnitude that we were measuring coming down the inlet during the duration of the flood.
It will be interesting to see what a more detailed dye mass/volume balance will show.
The attached video shows a 17 minute series of 10 second averaged images from the thermal IR cameras on the APLUW/OSU tower. Brighter shades indicate relatively warmer temperatures. The high tide was approximately 2 hours prior to the movie. Note the strong northward flow which bends into the inlet and intersects a cooler front in the main channel.
Here are the results of two AUV missions performed on 5/16 and 5/17 to observe the offshore extent of the river plume during max ebb using elevated CDOM signatures to define the boundaries of the plume. The 5/16 mission was performed during strong southerly wind conditions (blowing directly into the inlet) while the 5/17 mission was performed during minimal wind conditions. Both missions extend to 1km offshore. You’ll notice how defined the river jet is during the 5/17 mission and how the plume extends beyond our survey lines. In contrast, it appears the strong wind conditions during the 5/16 mission suppressed the offshore extent of the plume and pushed it in a southwesterly direction. It’s also interesting to see the impact the wind conditions have on the mixing of the plume. The high wind event on 5/16 diluted the plume significantly (colorbar range: 6 to 10 ppb) when compared to the low wind event on 5/17 where the river plume saw consistent CDOM levels of ~13 ppb as it is advected offshore.
Hello everyone, just wanted to pass along some of the work the SIO CORDC group has been performing offshore with the REMUS AUV. On 5/8/2012 we ran a REMUS mission supporting the Guza dye release. We’ve overlaid the Rhodamine signature observed by our AUV on the hyperspectral image taken at 17:39UTC (courtesy of Luc Lenain and Nick Statom). The REMUS was deployed at 17:20UTC and started near the entrance to the new channel and went in a southwesterly direction towards the beach. It reached the end of this leg at ~17:40UTC(when the image was taken). You can see from the image that this first leg was outside of the dye signature which corresponded with our observations The REMUS then performs a mow the law pattern across the inlet. We first encountered dye at ~17:46UTC and continued measuring the plume until the end of the mission at ~19:00UTC.